OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = x2 + 1, find f(g(3)).

OpenStudy (anonymous):

what is \(g(3)\) ?

OpenStudy (ivettef365):

first solve g(3), substitute the 3 for the x on function g(x)

OpenStudy (anonymous):

First evaluate g(3) by replacing the variable x with 3: \[g(3)=3^2+1\]\[g(3)=9+1=10\] So, now evaluate: \[f(10)=2*10+3=20+3=23\] Done.

OpenStudy (anonymous):

Thank you that helped alot i had kept forgeting to square it at first.

OpenStudy (anonymous):

No problem =]